# VBIM-Net: Variational Born Iterative Network for Inverse Scattering Problems

### v1

@@ IV-B System Configurations and Training Details
TABLE III: Performance Metrics of Reconstruction Results for Test#11-Test#20
@@
We mix the MNIST [42] and cylinder dataset at a ratio of 1:1 to obtain the synthetic dataset, in order to increase the richness of the data, making the network to learn inversion tasks for homogeneous and inhomogeneous scatterers with various shapes. For the cylinder dataset, each scene contains 1 to 3 cylinders with a 0.01m-0.05m radius. For lossless scatters, the contrast range is $[0.2,2.2]$. For lossy scatters, the range of the real part of the contrast is $[0.2,2.2]$, and the range of the imaginary part is $[0,1]$. Fig. 5 shows several examples of lossy scatters in the synthetic dataset. We randomly generate 20,000 data samples, of which 80% are for training and 20% for testing.
VBIM-Net and the modified NeuralBIM are implemented in PyTorch and computed on a Nvidia V100 GPU, and they share the same training configuration. Table I summarizes the implementation details of both. Note that since the modified NeuralBIM adds a trainable dimension transformation module to implement the mapping $\delta\mathbf{E}^{s}\rightarrow\mathbf{F}^{\delta\mathbf{E}^{s}}$, the number of parameters of the modified NeuralBIM is larger than that of VBIM-Net.
IV-C Synthetic Data Inversion: Within-Dataset Test
In this subsection, we first visually display the output of each VBIM-Net layer, and then test the model’s performance on the MNIST and cylinder datasets. Finally, we conduct an ablation experiment to verify the role of layer-wise constraint in the loss function.
IV-C1 Updated Contrast Reconstructions of VBIM-Net
IV-C2 Within-Dataset Samples Reconstruction
Fig. 7 shows several reconstruction results on the cylinder and MNIST test sets of these three algorithms when the noise level is 10%, and the corresponding performance metrics are summarized in Table II. Reconstruction results indicate that VBIM-Net outperforms SOM and NeuralBIM. Although both VBIM-Net and NeuralBIM are designed based on the field-type iterative method, the reconstruction quality of VBIM-Net is significantly better. This is mainly attributed to the analytical calculation of the approximate contrast variation $\delta\tilde{\boldsymbol{\chi}}$ that converts the scattered field residual $\delta\mathbf{E}^{s}$ into the contrast domain, which is more conducive to the subsequent CNN for image-to-image translation. Moreover, the U-Net used by VBIM-Net is more effective in extracting the multi-level features in contrast images than the ResNet in NeuralBIM, which also helps to improve the model performance.
IV-C3 Ablation Experiments for the Layer-Wise Constraint
In order to further verify the efficacy of the layer-wise constraint, we perform the following ablation experiment: set $c=0$ in Eq. (29) to only supervise the output of the last layer of VBIM-Net (where $10^{0}=$ 1 is defined), while keeping other settings unchanged. Fig. 9 shows the contrast MSE on the test set during training with/without the layer-wise constraint. When using the layer-wise constraint, the reconstruction quality gradually improves with the stacking of update blocks. With the same number of subnetwork layers, the model with the layer-wise constraint can converge to a better performance. Fig. 10 illustrates the reconstruction results without the layer-wise constraint, where the scatter is the same as in Fig. 6(a). It can be seen that removing the layer-wise constraint can result in an ambiguous physical interpretation of the output from intermediate layers, as well as a deterioration in the final reconstruction quality. This proves that the layer-wise constraint is necessary for VBIM-Net, which also validates the effectiveness of incorporating the physical constraint of iterative algorithms in network design.
@@ IV-D1 Cross-Dataset Samples Reconstruction
Fig. 11 shows the reconstruction results of Test#11-Test#20 with 10% Gaussian noise, and the reconstruction quality metrics are listed in Table III. It can be seen that VBIM-Net shows superior performance in cases where scatterers have complex shapes and high contrast. In the reconstruction of Test#17, divergence occurred in SOM, whereas VBIM-Net still achieves successful reconstruction. Test#19 tests the splicing of two scatterers with different contrasts, and Test#20 tests three dispersed scatterers with various shapes and contrasts. In terms of reconstruction quality of these test samples, VBIM-Net not only outperforms the iterative method SOM but also surpasses the modified NeuralBIM even with fewer learnable parameters. These results fully demonstrate the cross-dataset generalization ability of VBIM-Net, which is attributed to the physical information introduced in the network design and the training process.
@@
TABLE IV: Performance Metrics of Reconstruction Results for Test#21-Test#23
@@ IV-D2 Test under Different Noise Levels
We test the performance of VBIM-Net under different noise levels. We choose the “Austria” profile [43] with contrast $\chi=1.0$ as the target scatterer and test the reconstruction performance when adding 10% (SNR=20dB), 20% (SNR=14dB), and 30% (SNR=10dB) Gaussian noise. The reconstruction results of Test#21-Test#23 under different noise levels are shown in Fig. 12, and their corresponding performance metrics are listed in Table IV. Experimental results show that VBIM-Net can achieve better reconstruction quality than other two methods under various noise levels. This is mainly due to VBIM-Net’s physics-inspired iterative framework and the training scheme with extra noise, which enhance the network’s robustness.
@@ IV-D4 Test with Different Measurement Dimensions
TABLE V: Performance Metrics of the Real Part of the Reconstruction Results for Test#29-Test#35
@@
TABLE VI: Performance Metrics of the Imaginary Part of the Reconstruction Results for Test#29-Test#35
@@
IV-D5 Test with Lossy Scatterers
Finally, we test VBIM-Net’s reconstruction ability for lossy scatterers. We use the same MNIST-cylinder hybrid synthetic training dataset configuration as the lossless case. The only difference is that all scatterers have an imaginary part between 0 and 1.0. For the VBIM-Net model, we only need to modify the input and output channel number of contrast variation update network $\mathcal{F}^{\delta\boldsymbol{\chi}}_{k}$, keeping other model parameters and hyperparameters unchanged. The real and imaginary parts of the reconstructed contrast for Test#29-Test#35 are shown in Fig. 18, and their corresponding quality metrics are summarized in Tables V and VI. It demonstrates that VBIM-Net can perform high-quality reconstruction of diverse lossy scatterers. Compared with the lossless situation, VBIM-Net exhibits more remarkable performance advantages than traditional iterative algorithms such as SOM. These experimental results further verify the effectiveness of VBIM-Net in addressing ISPs.
@@ IV-E Experimental Data Inversion
In this section, we test VBIM-Net on experimental data from the Fresnel Institute [44]. We considered the “FoamDielExt” profile and the “FoamDielInt” profile, as shown in Fig. 19. They consist of a foam cylinder with a diameter of 80mm and a relative permittivity $\varepsilon_{r}^{\mathrm{f}}=1.45\pm 0.15$, and a plastic cylinder with a diameter of 31mm and a relative permittivity $\varepsilon_{r}^{\mathrm{p}}=3.0\pm 0.3$. Different from the previous configuration of synthetic data, there are 8 TXs and 241 RXs located on a circle with a radius of 1.67m, and the RXs are evenly distributed in the angular range of $60^{\circ}$-$300^{\circ}$ relative to each TX. We select the measurement data with operating frequencies of 3 GHz and 4 GHz, keeping the DOI size at $0.2\mathrm{m}\times 0.2\mathrm{m}$ and the inversion grid resolution at $64\times 64$. Although the experimental scene only contains cylindrical scatterers, we still synthesize the training dataset by mixing MNIST and cylinder scatterers, with the purpose of reducing the reliance on the dataset prior and avoiding the loss of generalization ability of the model. The hyperparameter $c$ in the layer-wise constraint is adjusted to 0.7, while the rest of the training configuration remains unchanged.
@@
TABLE VII: Performance Metrics of the Reconstruction Results of “FoamDielExt” and “FoamDielInt” Profiles at 3 GHz and 4 GHz
@@
SOM 0.81 0.16 0.80 0.21 0.74 0.16 0.73 0.17
NeuralBIM 0.89 0.07 0.87 0.17 0.74 0.32 0.86 0.20
VBIM-Net 0.98 0.04 0.98 0.05 0.94 0.07 0.90 0.16
(a) $f=3\,\mathrm{GHz}$.
(b) $f=4\,\mathrm{GHz}$.
Fig. 20 shows the reconstruction results of “FoamDielExt” and “FoamDielInt” profiles at 3 GHz and 4 GHz, and the corresponding quality metrics are summarized in Table VII. Fig. 21 shows the predicted scattered field of VBIM-Net for the “FoamDielExt” profile. The results based on experimental data demonstrate that VBIM-Net achieves preferable contrast reconstructions and accurate scattering field predictions at different frequencies, which further verify the effectiveness of the proposed VBIM-Net.
@@ V Conclusion
In this paper, VBIM-Net is proposed to solve full-wave ISPs. VBIM-Net emulates the iterative process of VBIM by multiple layers of subnetworks. In the contrast update step, we convert the scattered field residual into an approximate contrast variation based on VBIM and then enhance it by a U-Net, which avoids the requirement for matched measurement dimension and grid resolution in NeuralBIM. To guarantee the consistency between predicted physical quantities and their actual value in each subnetwork, we introduce a layer-wise constraint in the loss function, which makes the data flow of VBIM-Net comply with an iterative process, thereby ensuring the model’s reliability. In the training stage, we enhance the model’s generalization ability by mixing the cylinder and MNIST datasets, and design a noisy training scheme to improve the model’s robustness. Test results on synthetic and experimental data verify the superior performance of VBIM-Net compared to SOM and NeuralBIM. Overall, VBIM-Net has high reconstruction quality, strong generalization ability, and good physical interpretability. It will provide new ideas for the design of field-type DL schemes.
Despite the improvement achieved, the proposed VBIM-Net still needs to be further investigated in some aspects, such as semi-supervised and unsupervised learning schemes and decoupling the model from the measurement settings.

### v2

@@ IV-C4 Validation of the Efficacy of the Layer-Wise Constraint
Figure 12: The contrast SSIM performance curve of VBIM-Net with different layer numbers $K$, trained under different layer-wise constraint parameters $c$. The performance is evaluated on the test set.
@@
TABLE IV: Performance Metrics of Reconstruction Results for Test#11-Test#20
@@
To further reveal the effect of the layer-wise constraint parameter on the performance of VBIM-Net, we plot the contrast SSIM curve of VBIM-Net with different $K$ and $c$ on the test set in Fig. 12. When $K=5$, the network has too few layers and lacks the iterative characteristics typically expected. In this case, the impact of $c$ on model performance is not clearly patterned. As the number of network layers increases to $K~{}=~{}7,~{}9,~{}11$, the effect of the layer-wise constraint becomes more evident, and the optimal value for parameter $c$ shows a gradual increase. The above observations suggest that the layer-wise constraint is critical for training deep VBIM-Net models. In the absence of the layer-wise constraint ($c=0$), the performance of the model may deteriorate as the model depth $K$ increases. This is attributed to the overfitting caused by the lack of physical constraints. Nevertheless, the incorporation of the layer-wise constraint can enhance the stability during training for deep VBIM-Net models, enabling a higher performance upper bound with more update layers. Although there is a correlation between the optimal value for parameter $c$ and the number of network layers $K$, in practice, setting $c$ as the default value of 0.8 can achieve satisfactory performance in most cases.
Since the layer-wise constraint requires intermediate layer outputs and increases the complexity of the loss function, it indeed slightly raises the computational cost of optimizing VBIM-Net. According to the experimental results, with the same training configurations, adding the layer-wise constraint increases the training time by $2.5\%$, but it effectively enhances the model’s interpretability and scalability.
The above experiments demonstrate the efficacy of the layer-wise constraint in training deep-unfolding models like VBIM-Net, especially when we need to enhance the model performance through increasing the model’s depth. This also validates the effectiveness of incorporating soft physical constraints from iterative algorithms into the network design.
@@ IV-D1 Cross-Dataset Samples Reconstruction
Fig. 13 shows the reconstruction results of Test#11-Test#20 with 10% Gaussian noise, and the reconstruction quality metrics are listed in Table IV. It can be seen that VBIM-Net shows superior performance in cases where scatterers have complex shapes and high contrast. In the reconstruction of Test#17, divergence occurred in BIM, VBIM and SOM, whereas VBIM-Net still achieves successful reconstruction. Test#19 tests the splicing of two scatterers with different contrasts, and Test#20 tests three dispersed scatterers with various shapes and contrasts. In terms of reconstruction quality of these test samples, VBIM-Net not only outperforms iterative methods such as BIM, VBIM and SOM, but also surpasses the modified NeuralBIM, even with fewer learnable parameters. These results fully demonstrate the cross-dataset generalization ability of VBIM-Net, which is attributed to the physical information introduced in the network architecture and the loss function. Given the significant performance disadvantage of BIM and VBIM in both within-dataset and cross-dataset tests, we exclude them from the comparisons in the following experiments.
@@
TABLE V: Performance Metrics of Reconstruction Results for Test#21-Test#23
@@ IV-D2 Test under Different Noise Levels
We test the performance of VBIM-Net under different noise levels. We choose the “Austria” profile [43] with contrast $\chi=1.0$ as the target scatterer and test the reconstruction performance when adding 10% (SNR=20dB), 20% (SNR=14dB), and 30% (SNR=10dB) Gaussian noise. The reconstruction results of Test#21-Test#23 under different noise levels are shown in Fig. 14, and their corresponding performance metrics are listed in Table V. Experimental results show that VBIM-Net can achieve better reconstruction quality than other two methods under various noise levels. This is mainly due to VBIM-Net’s physics-inspired iterative framework and the training scheme with extra noise, which enhance the network’s robustness.
@@ IV-D3 Test with Different Contrast
TABLE VI: Performance Metrics of the Real Part of the Reconstruction Results for Test#29-Test#35
@@
TABLE VII: Performance Metrics of the Imaginary Part of the Reconstruction Results for Test#29-Test#35
@@
IV-D4 Test with Different Measurement Dimensions
IV-D5 Test with Lossy Scatterers
Finally, we test VBIM-Net’s reconstruction ability for lossy scatterers. For the VBIM-Net model, we only need to modify the input and output channel number of contrast variation update network $\mathcal{F}^{\delta\boldsymbol{\chi}}_{k}$, keeping other model parameters and hyperparameters unchanged. We switch to using the lossy scatterer dataset for training models. The real and imaginary parts of the reconstructed contrast for Test#29-Test#35 are shown in Fig. 20, and their corresponding quality metrics are summarized in Tables VI and VII. It demonstrates that VBIM-Net can perform high-quality reconstruction of diverse lossy scatterers. Compared with the lossless situation, VBIM-Net exhibits more remarkable performance advantages than traditional iterative algorithms such as SOM. These experimental results further verify the effectiveness of VBIM-Net in addressing ISPs.
@@ IV-E Experimental Data Inversion
In this section, we test VBIM-Net on experimental data from the Fresnel Institute [44]. We considered the “FoamDielExt” profile and the “FoamDielInt” profile, as shown in Fig. 21. They consist of a foam cylinder with a diameter of 80mm and a relative permittivity $\varepsilon_{r}^{\mathrm{f}}=1.45\pm 0.15$, and a plastic cylinder with a diameter of 31mm and a relative permittivity $\varepsilon_{r}^{\mathrm{p}}=3.0\pm 0.3$. Different from the previous configuration of synthetic data, there are 8 TXs and 241 RXs located on a circle with a radius of 1.67m, and the RXs are evenly distributed in the angular range of $60^{\circ}$-$300^{\circ}$ relative to each TX. We select the measurement data with operating frequencies of 3 GHz and 4 GHz, keeping the DOI size at $0.2\mathrm{m}\times 0.2\mathrm{m}$ and the inversion grid resolution at $64\times 64$. We use the lossless dataset with cylindrical scatterers mentioned in Section IV-B as training data. The model configurations are consistent with the default settings of the synthetic data inversion experiments.
Fig. 22 shows the reconstruction results of “FoamDielExt” and “FoamDielInt” profiles at 3 GHz and 4 GHz, and the corresponding quality metrics are summarized in Table VIII. Fig. 23 shows the predicted scattered field of VBIM-Net for the “FoamDielExt” profile. The results based on experimental data demonstrate that VBIM-Net achieves preferable contrast reconstructions and accurate scattered field predictions at different frequencies, which further verify the effectiveness of the proposed VBIM-Net.
@@
(a) $f=3\,\mathrm{GHz}$.
(b) $f=4\,\mathrm{GHz}$.
TABLE VIII: Performance Metrics of the Reconstruction Results of “FoamDielExt” and “FoamDielInt” Profiles at 3 GHz and 4 GHz
@@
SOM 0.81 0.16 0.80 0.21 0.78 0.13 0.81 0.14
NeuralBIM 0.94 0.04 0.88 0.12 0.76 0.26 0.80 0.19
VBIM-Net 0.98 0.04 0.98 0.05 0.95 0.06 0.95 0.10
@@ V Conclusion
In this paper, we propose VBIM-Net, a field-type DL scheme designed to solve full-wave ISPs by emulating the iterative process of VBIM through multiple layers of subnetworks. Inspired by the VBIM algorithm, we incorporate an analytical computation in the contrast update step, which converts the scattered field residual into an approximate contrast variation and then refines it using a U-Net. This approach avoids the need for matched measurement dimensions and grid resolution required in NeuralBIM. To ensure the consistency between predicted and actual physical quantities within each subnetwork, we introduce a layer-wise constraint in the loss function, which serves as a soft physical constraint. This constraint enforces the data flow of VBIM-Net comply with an iterative process, thereby enhancing the reliability and generalization of deep models. In the training stage, we adopt a noisy training scheme to further improve the model’s robustness. Testing results on synthetic and experimental data verify the superior performance of VBIM-Net compared to BIM, SOM and NeuralBIM, and a series of experiments also prove the significant effect of the new design we introduced. Overall, VBIM-Net achieves high reconstruction quality, strong generalization ability, and good physical interpretability, providing new ideas for the design of field-type DL schemes. Despite these advancements, the proposed VBIM-Net still can be further explored in some aspects, such as semi-supervised and unsupervised learning schemes, fully decoupling the model from the measurement settings, and flexible reconstruction resolutions.